Grade 12th passMechanics

A body of mass 4kg has an initial velocity of 6m/s along x axis and it is subjected to a force of 8N in y direction .the displavement of the body from origin after 2sec will be (the body was subjected to force at origin at t=0)
1) 12.65m
2)16m
3)14m
4)18m

sheema
9 Years agoGrade 12th pass

1 Answer

Vaibhav Singh
9 Years ago
(A) is correct.Since the body is moving with constant velocity of 6 m/s in x direction (because there is no force in the x direction). In y direction there is no initial velocity and a force of 8 N start acting. Therefore acceleration in y direction will be 8N/4Kg = 2 m/s^2.Now displacement in y direction will be Sy = 0(2) +1/2(2)2^2. (Sy =ut + 1/2at^2, here u is 0).Similarly, displacement in x direction will be Sx = 6×2 (ie Sx=v×t).Therefore Sy= 4, Sx= 12.Total displacement S = √(Sx^2 +Sy^2)= √(12^2 + 4^2) ~= 12.7.Since, 12.65 is closest it will be the correct answer.